# From Freshness to Effectiveness: Goal-Oriented Sampling for Remote Decision Making

### v2

@@ V-B Primal-Dinkelbach Synchronous Iteration (Proposed OnePDSI)
(OnePDSI): For a given $10<\kappa<$ 1, we can iteratively generate sequences $\{\rho_{K}\}^{K\in\mathbb{N}^{+}}$ and $\{\widetilde{W}_{K}(\gamma)\}_{\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}^{K\in\mathbb{N}^{+}}$ with a starting initial value $\{\widetilde{W}_{0}(\gamma)\}_{\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}$:
$\rho_{K+1}$ $=\min_{A_{i},Z_{i}}\Bigg{\{}\frac{q(\gamma^{\text{r}},Z_{i},A_{i})-\kappa\widetilde{W}_{K}(\gamma^{\text{r}})\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}+\frac{\kappa\mathbb{E}\left[\widetilde{W}_{K}(\gamma^{\prime})\mid\gamma^{\text{r}},Z_{i},A_{i}\right]\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}\Bigg{\}}+\widetilde{W}_{K}(\gamma^{\text{r}}),$ (39a)
$\widetilde{W}_{K+1}(\gamma)$ $=\min_{A_{i},Z_{i}}\Bigg{\{}\frac{q(\gamma,Z_{i},A_{i})-\kappa\widetilde{W}_{K}(\gamma)\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}+\frac{\kappa\mathbb{E}\left[\widetilde{W}_{K}(\gamma^{\prime})\mid\gamma,Z_{i},A_{i}\right]\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}\Bigg{\}}+\widetilde{W}_{K}(\gamma)-\rho_{K+1},\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}.$ (39b)
@@ V-C Convergence of OnePDSI
(Convergence of OnePDSI). If the transformed MDP is unichain and $10<\kappa<$ 1, then the OnePDSI in (39) is convergent, with:
@@
See Appendix LABEL:appendixJ. ∎
The next theorem characterizes an upper bound on the relative error of the proposed OnePDSI, whose proof is provided in Appendix LABEL:appendixJ:
@@
(Upper Bound of Relative Error). If the MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ is a unichain MDP, then up to iteration $K$, the relative error $e_{\mathrm{\rho}}^{(K)}(\gamma)$ is upper bounded above by
where $M$ is a scaling factor, $L$ is defined by (LABEL:181). The term $R=\frac{1}{(1-\epsilon)^{1/L}}$ captures the asymptotic convergence rate.
@@ VI-A Lagrangian Dual Techniques
Following the steps in (11)-(14) and applying Dinkelbach’s method for non-linear fractional programming as in [89] and [20, Lemma 2], the problem of determining the optimal policy for Problem 1 is equivalent to solving the following alternative problem given the Dinkelbach parameter $\lambda$:
@@
$H(\lambda;f_{\max})\triangleq$ $\inf_{\boldsymbol{\psi}}\lim_{\mathrm{n}\rightarrow\infty}\frac{1}{n}{\sum_{i=0}^{n-1}\left\{\mathbb{E}_{\boldsymbol{\psi}}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]-\lambda\mathbb{E}_{\boldsymbol{\psi}}\left[Z_{i}+Y_{i+1}\right]\right\}}$ (44)
[…]
The following lemma characterizes the relationship between $h^{\star}$ and the optimal value of Problem 5. The proof follows directly from modifying the policies in [1, Appendix B] to satisfy the sampling frequency constraint, and is therefore omitted.
@@
(ii). When $H(\lambda;f_{\max})=0$, the solution to Problem 5 coincide with those of Problem Problem 1.
@@
With Lemma 6 in hand, solving Problem 1 is equivalent to solving for the root of the implicit function $H(\lambda;f_{\max})$. To obtain $H(\lambda;f_{\max})$ given $\lambda$, we first transform the CMDP into an unconstrained Lagrangian MDP problem. Specifically, define the Lagrange Dual Function as:
$\mathcal{L}(\boldsymbol{\psi};\theta,\lambda,f_{\max})=\frac{\theta}{f_{\max}}+\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\boldsymbol{\psi}}\left\{\sum_{t=D_{i}}^{D_{i+1}-1}\mathcal{C}\left(X_{t},A_{i}\right)-(\lambda+\theta)\left(Z_{i}+Y_{i+1}\right)\right\},$ (45)
@@
where $\Upsilon(\theta,\lambda;f_{\max})=\inf_{\phi}\mathcal{L}(\phi;\theta,\lambda,f_{\max})$
The weak duality principle [95, Chapter 5.2.2] implies that $d(\lambda;f_{\max})$ is a lower bound of $H(\lambda;f_{\max})$, i.e., $d(\lambda;f_{\max})\leq H(\lambda;f_{\max})$. In the following lemma, we establish the conditions where the strong duality holds true and thus $d(\lambda;f_{\max})=H(\lambda;f_{\max})$. Under these conditions, it is sufficient to solve $d(\lambda;f_{\max})$ to obtain $H(\lambda;f_{\max})$.
@@
(Restatement of [95, Chapter 5.5.3]) The duality gap between Problem 5 and Problem 6 is zero, i.e., $d(\lambda;f_{\max})=H(\lambda;f_{\max})$, if and only if for any given $\lambda$, we can find $\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}$ and $\theta^{\star}_{\lambda}$ such that the Karush–Kuhn–Tucker (KKT) conditions are satisfied:
@@
$\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}=\mathop{\arg\min}_{\phi}\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\phi}\Bigg{\{}\sum_{t=D_{i}}^{D_{i+1}-1}\mathcal{C}\left(X_{t},A_{i}\right)-(\lambda+\theta^{\star}_{\lambda})\left(Z_{i}+Y_{i+1}\right)\Bigg{\}}+\frac{\theta^{\star}_{\lambda}}{f_{\max}},$ (48b)
@@ VI-B Three-Layer Solutions and the Structure of Optimal Policies
In this subsection, we propose a three-layer algorithm outlined in Algorithm 2. The basic framework of this algorithm is inspired by [85, Section IV.C]. This algorithm consists of inner, middle, and outer layers, whose implementations are detailed below.
@@ VI-B1 Inner-Layer: A Standard MDP Given $\lambda$ and $\theta$
For any given $\theta$ and $\lambda$, the inner layer $\inf_{\phi}\mathcal{L}(\phi;\theta,\lambda,f_{\max})$ is a standard unconstrained infinite horizon MDP as defined in IV-C1, which is denoted by $\mathscr{P}_{\text{MDP}}(\theta+\lambda)$ with the optimal value $U(\lambda+\theta)$. This problem can be efficiently solved using the $\tau$-RVI algorithm, as outlined in Iteration 2. Notably, [85, Section IV.C] applies a value iteration process in the inner layer, while we apply the $\tau$-RVI to ensure the rigorous convergence in our context.
@@ VI-B2 Middle-Layer: Update Lagrangian multiplier $\theta$ Given $\lambda$
Given a Dinkelbach parameter $\lambda$, the middle layer involves solving $\max_{\theta\geq 0}\Upsilon(\theta,\lambda;f_{\max})$ in Problem 6 to accurately approximate $d(\lambda;f_{\max})$, which searches for the optimal $\theta^{\star}_{\lambda}$ and its corresponding optimal policy $\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}$ that satisfy the KKT conditions in Lemma 7. Different from [85, Section IV.C] where the sub-gradient method is employed to update the Lagrangian multiplier leveraging the convexity of the Lagrangian Dual function, we conduct a monotonicity analysis and explicitly state the conditions for the optimal multiplier. We begin by introducing the following notations that help clarify the role of the multiplier $\theta$ given a fixed $\lambda$. Specifically, denote
@@
where $\phi^{\star}_{\lambda+\theta}$ is the optimal stationary deterministic policy for the unconstrained problem $\mathscr{P}_{\text{MDP}}(\lambda+\theta)$ given in IV-C1. The subsequent Lemma presents key properties of $\mathcal{F}^{\lambda+\theta}$, $\mathcal{Q}^{\lambda+\theta}$, $U(\lambda+\theta)$, and $\Upsilon(\theta,\lambda;f_{\max})$ with respect to $\theta$, which are useful for the explicit solutions.
@@
(A variant of [96, Lemma 3.1]) The following assertions hold true:
($i$) $U({\lambda+\theta})$ is non-increasing with respect to $\theta$;
($ii$) $\mathcal{Q}^{\lambda+\theta}$ and $\mathcal{F}^{\lambda+\theta}$ are non-decreasing functions with respect to $\theta$;
($iii$) $\mathcal{Q}^{\lambda+\theta}$ and $\mathcal{F}^{\lambda+\theta}$ are both step functions with respect to $\theta$;
($iv$) If $\mathcal{F}^{\lambda+\theta}\geq 1/f_{\max}$, $\Upsilon(\theta,\lambda;f_{\max})$ is non-increasing with respect to $\theta$.
@@
Part ($i$) and part ($ii$) are supported by [96, Lemma 3.1]. Part ($iii$) is supported by [96, Lemma 3.2]. See Appendix LABEL:appendixkv3 for the detailed proof of part ($iv$). ∎
@@
Figure 6: Illustrations of Case ($i$) and case ($ii$) in Corollary 1. In this figure, $\mathcal{F}^{\lambda+\theta}$ is a non-increasing step function with respect to $\theta$, as Lemma 8-($ii$) and Lemma 8-($iii$) indicate. In addition, $\Upsilon(\theta,\lambda;f_{\max})$ is non-increasing with respect to $\theta$ if $\mathcal{F^{\lambda+\theta}}\geq 1/f_{\max}$, as Lemma 8-($iv$) indicates.
@@
If $\mathcal{F}^{\lambda+}\geq\frac{1}{f_{\max}}$, which is illustrated in the right panel of Fig. 6, the function $\Upsilon(\theta,\lambda;f_{\max})$ is non-increasing with $\theta$ for $\theta\geq 0$. In this case, the maximum value of $\Upsilon(\theta,\lambda;f_{\max})$ is obtained at $\theta_{\lambda}^{\star}=0$. If $\mathcal{F}^{\lambda+}<\frac{1}{f_{\max}}$, the feasible region under the KKT condition (48c) is shown in the left panel of Fig. 6 with $\theta\geq\theta^{\text{tr}}$. In this feasible region, the function $\Upsilon(\theta,\lambda;f_{\max})$ is non-increasing in $\theta$. Therefore, the maximum value of $\Upsilon(\theta,\lambda;f_{\max})$ is obtained at $\theta_{\lambda}^{\star}=\theta^{\text{tr}}$. A detailed proof is provided in Appendix LABEL:appendixlv2. ∎
Having established Corollary 1, our remaining task is to identify the threshold value $\theta^{\text{tr}}$ that satisfies (52) under the condition that $\mathcal{F}^{\lambda^{+}}<1/f_{\max}$. Given that $\mathcal{F}^{\lambda+\theta}$ is non-decreasing in $\theta$, we can apply a bisection search to gradually converge on the threshold $\theta^{\text{tr}}$. Once this threshold value is obtained, setting $\theta_{\lambda}^{\star}=\theta^{\text{tr}}$ ensures that the KKT conditions (48a)–(48c) are satisfied. What remains is to determine the optimal policy $\phi_{\lambda+\theta^{\star}_{\lambda}}$ that guarantees the KKT condition (48d). The structure of this optimal policy is provided in the following theorem.
@@
($i$). If $\mathcal{F}^{\lambda^{+}}\geq 1/f_{\max}$, the optimal policy is a stationary deterministic policy $\phi^{\star}_{\lambda}$ determined in (IV-C1);
($ii$). If $\mathcal{F}^{\lambda^{+}}<1/f_{\max}$ and $\mathcal{F}^{(\lambda+\theta^{\star}_{\lambda})^{+}}=1/f_{\max}$, the optimal policy is a stationary deterministic policy $\phi^{\star}_{(\lambda+\theta^{\star}_{\lambda})}$;
($iii$). If $\mathcal{F}^{\lambda^{+}}<1/f_{\max}$ and $\mathcal{F}^{(\lambda+\theta^{\star}_{\lambda})^{+}}>1/f_{\max}$, the optimal policy is a random mixture of two stationary deterministic policy $\phi^{\star}_{(\lambda+\theta^{\star}_{\lambda})^{+}}$ and $\phi^{\star}_{(\lambda+\theta^{\star}_{\lambda})^{-}}$:
@@
=$=\begin{cases}U(\lambda)&\text{if }\mathcal{F}^{\lambda^{+}}\geq\frac{1}{f_{\max}}\\ U(\lambda+\theta_{\lambda}^{\star})+\frac{\theta_{\lambda}^{\star}}{f_{\max}}&\text{if }\mathcal{F}^{\lambda^{+}}<\frac{1}{f_{\max}}\end{cases}.$

### v3

@@ V-B Primal-Dinkelbach Synchronous Iteration (Proposed OnePDSI)
(OnePDSI): For a given $0<\kappa<1$, we can iteratively generate sequences $\{\rho_{K}\}^{K\in\mathbb{N}^{+}}$ and $\{\widetilde{W}_{K}(\gamma)\}_{\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}^{K\in\mathbb{N}^{+}}$ with a starting initial value $\{\widetilde{W}_{0}(\gamma)\}_{\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}$:
$\rho_{K+1}$ $=\min_{A_{i},Z_{i}}\Bigg\{\frac{q(\gamma^{\text{r}},Z_{i},A_{i})-\kappa\widetilde{W}_{K}(\gamma^{\text{r}})\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}$
$+\frac{\kappa\mathbb{E}\left[\widetilde{W}_{K}(\gamma^{\prime})\mid\gamma^{\text{r}},Z_{i},A_{i}\right]\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}\Bigg\}+\widetilde{W}_{K}(\gamma^{\text{r}}),$ (43a)
$\widetilde{W}_{K+1}(\gamma)$ $=\min_{A_{i},Z_{i}}\Bigg\{\frac{q(\gamma,Z_{i},A_{i})-\kappa\widetilde{W}_{K}(\gamma)\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}$
$+\frac{\kappa\mathbb{E}\left[\widetilde{W}_{K}(\gamma^{\prime})\mid\gamma,Z_{i},A_{i}\right]\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}\Bigg\}$
$+\widetilde{W}_{K}(\gamma)-\rho_{K+1},\quad\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}.$ (43b)
@@ V-C Convergence of OnePDSI
(Convergence of OnePDSI). If the transformed MDP is unichain and $0<\kappa<1$, then the OnePDSI in (43) is convergent, with:
@@
See Appendix J. ∎
The next theorem characterizes an upper bound on the relative error of the proposed OnePDSI, whose proof is provided in Appendix J:
@@
(Upper Bound of Relative Error). If the MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ is a unichain MDP, then up to iteration $K$, the relative error $e_{\mathrm{\rho}}^{(K)}$ is upper bounded above by
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}e_{\rho}^{(K)}\leq\frac{\tau M(1-\epsilon)^{(K-1)/L}}{1-(1-\epsilon)^{1/L}}=\mathcal{O}\left(\frac{1}{R^{K}}\right),$ (47)
where $M$ is a scaling factor, $L$ is defined by (182). The term $R=\frac{1}{(1-\epsilon)^{1/L}}$ captures the asymptotic convergence rate.
@@ VI-A Lagrangian Dual Techniques
Following the steps in (14a)-(17) and applying Dinkelbach’s method for non-linear fractional programming as in [28] and [85, Lemma 2], the problem of determining the optimal policy for Problem 1 is equivalent to solving the following alternative problem given the Dinkelbach parameter $\lambda$:
@@
$H(\lambda;f_{\max})\triangleq$ (48)
$\inf_{\boldsymbol{\psi}}\lim_{\mathrm{n}\rightarrow\infty}\frac{1}{n}{\sum_{i=0}^{n-1}\left\{\mathbb{E}_{\boldsymbol{\psi}}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]-\lambda\mathbb{E}_{\boldsymbol{\psi}}\left[Z_{i}+Y_{i+1}\right]\right\}}$
$\text{s.t. }\lim_{T\to\infty}\frac{1}{T}\mathbb{E}_{\boldsymbol{\psi}}\left[\sum_{i=1}^{T}(S_{i+1}-S_{i})\right]\geq\frac{1}{f_{\text{max}}},$
The following lemma characterizes the relationship between $h^{\star}$ and the optimal value of Problem 5. The proof follows directly from modifying the policies in [49, Appendix B] to satisfy the sampling frequency constraint, and is therefore omitted.
@@
(ii). When $H(\lambda;f_{\max})=0$, the solutions to Problem 5 coincide with those of Problem 1.
@@
With Lemma 6 in hand, solving Problem 1 is equivalent to solving for the root of the implicit function $H(\lambda;f_{\max})$. To obtain $H(\lambda;f_{\max})$ given $\lambda$, we first transform the CMDP into an unconstrained Lagrangian MDP problem. Specifically, define the Lagrange Function as:
$\mathcal{L}(\boldsymbol{\psi};\theta,\lambda,f_{\max})=\frac{\theta}{f_{\max}}+$ (49)
$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\boldsymbol{\psi}}\left\{\sum_{t=D_{i}}^{D_{i+1}-1}\mathcal{C}\left(X_{t},A_{i}\right)-(\lambda+\theta)\left(Z_{i}+Y_{i+1}\right)\right\},$
@@
where $\Upsilon(\theta,\lambda;f_{\max})=\inf_{\phi}\mathcal{L}(\phi;\theta,\lambda,f_{\max})$ .
The weak duality principle [15, Chapter 5.2.2] implies that $d(\lambda;f_{\max})$ is a lower bound of $H(\lambda;f_{\max})$, i.e., $d(\lambda;f_{\max})\leq H(\lambda;f_{\max})$. In the following lemma, we establish the conditions where the strong duality holds true and thus $d(\lambda;f_{\max})=H(\lambda;f_{\max})$. Under these conditions, it is sufficient to solve $d(\lambda;f_{\max})$ to obtain $H(\lambda;f_{\max})$.
@@
(Restatement of [15, Chapter 5.5.3]) The duality gap between Problem 5 and Problem 6 is zero, i.e., $d(\lambda;f_{\max})=H(\lambda;f_{\max})$, if and only if for any given $\lambda$, we can find $\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}$ and $\theta^{\star}_{\lambda}$ such that the Karush–Kuhn–Tucker (KKT) conditions are satisfied:
@@
$\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}=\mathop{\arg\min}_{\phi}\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\phi}\Bigg\{\sum_{t=D_{i}}^{D_{i+1}-1}\mathcal{C}\left(X_{t},A_{i}\right)$-
$\quad\quad\quad\quad\quad\quad\quad\quad\quad(\lambda+\theta^{\star}_{\lambda})\left(Z_{i}+Y_{i+1}\right)\Bigg\}+\frac{\theta^{\star}_{\lambda}}{f_{\max}},$ (52b)
@@ VI-B Three-Layer Solutions and the Structure of Optimal Policies
In this subsection, we propose a three-layer algorithm outlined in Algorithm 2. The basic framework of this algorithm is inspired by [110, Section IV.C]. This algorithm consists of inner, middle, and outer layers, whose implementations are detailed below.
@@ VI-B1 Inner-Layer: A Standard MDP Given $\lambda$ and $\theta$
For any given $\theta$ and $\lambda$, the inner layer $\inf_{\phi}\mathcal{L}(\phi;\theta,\lambda,f_{\max})$ is a standard unconstrained infinite horizon MDP as defined in IV-B1, which is denoted by $\mathscr{P}_{\text{MDP}}(\theta+\lambda)$ with the optimal value $U(\lambda+\theta)$. This problem can be efficiently solved using the $\tau$-RVI algorithm, as outlined in Iteration 2. Notably, [110, Section IV.C] applies a value iteration process in the inner layer, while we apply the $\tau$-RVI to ensure the rigorous convergence in our context.
@@ VI-B2 Middle-Layer: Update Lagrangian multiplier $\theta$ Given $\lambda$
Given a Dinkelbach parameter $\lambda$, the middle layer involves solving $\max_{\theta\geq 0}\Upsilon(\theta,\lambda;f_{\max})$ in Problem 6 to accurately approximate $d(\lambda;f_{\max})$, which searches for the optimal $\theta^{\star}_{\lambda}$ and its corresponding optimal policy $\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}$ that satisfy the KKT conditions in Lemma 7. Different from [110, Section IV.C] where the sub-gradient method is employed to update the Lagrangian multiplier leveraging the convexity of the Lagrangian Dual function, we conduct a monotonicity analysis and explicitly state the conditions for the optimal multiplier. We begin by introducing the following notations that help clarify the role of the multiplier $\theta$ given a fixed $\lambda$. Specifically,
@@
where $\phi^{\star}_{\lambda+\theta}$ is the optimal stationary deterministic policy for the unconstrained problem $\mathscr{P}_{\text{MDP}}(\lambda+\theta)$ given in IV-B1. The subsequent Lemma presents key properties of $\mathcal{F}^{\lambda+\theta}$, $\mathcal{Q}^{\lambda+\theta}$, $U(\lambda+\theta)$, and $\Upsilon(\theta,\lambda;f_{\max})$ with respect to $\theta$, which are useful for the explicit solutions.
@@
(A variant of [14, Lemma 3.1]) The following assertions hold true:
(ii) $U({\lambda+\theta})$ is non-increasing with respect to $\theta$;
(iiii) $\mathcal{Q}^{\lambda+\theta}$ and $\mathcal{F}^{\lambda+\theta}$ are non-decreasing functions with respect to $\theta$;
(iiiiii) $\mathcal{Q}^{\lambda+\theta}$ and $\mathcal{F}^{\lambda+\theta}$ are both step functions with respect to $\theta$;
(iviv) If $\mathcal{F}^{\lambda+\theta}\geq 1/f_{\max}$, $\Upsilon(\theta,\lambda;f_{\max})$ is non-increasing with respect to $\theta$.
@@
Part (ii) and part ($ii$) are supported by [14, Lemma 3.1]. Part (iiiiii) is supported by [14, Lemma 3.2]. See Appendix K for the detailed proof of part ($iv$). ∎
@@
Figure 6: Illustrations of Case (ii) and case (iiii) in Corollary 1. In this figure, $\mathcal{F}^{\lambda+\theta}$ is a non-increasing step function with respect to $\theta$, as Lemma 8-(iiii) and Lemma 8-(iiiiii) indicate. In addition, $\Upsilon(\theta,\lambda;f_{\max})$ is non-increasing with respect to $\theta$ if $\mathcal{F^{\lambda+\theta}}\geq 1/f_{\max}$, as Lemma 8-(iviv) indicates.
@@
If $\mathcal{F}^{\lambda+}\geq\frac{1}{f_{\max}}$, which is illustrated in the right panel of Fig. 6, the function $\Upsilon(\theta,\lambda;f_{\max})$ is non-increasing with $\theta$ for $\theta\geq 0$. In this case, the maximum value of $\Upsilon(\theta,\lambda;f_{\max})$ is obtained at $\theta_{\lambda}^{\star}=0$. If $\mathcal{F}^{\lambda+}<\frac{1}{f_{\max}}$, the feasible region under the KKT condition (52c) is shown in the left panel of Fig. 6 with $\theta\geq\theta^{\text{tr}}$. In this feasible region, the function $\Upsilon(\theta,\lambda;f_{\max})$ is non-increasing in $\theta$. Therefore, the maximum value of $\Upsilon(\theta,\lambda;f_{\max})$ is obtained at $\theta_{\lambda}^{\star}=\theta^{\text{tr}}$. A detailed proof is provided in Appendix L. ∎
Having established Corollary 1, our remaining task is to identify the threshold value $\theta^{\text{tr}}$ that satisfies (56) under the condition that $\mathcal{F}^{\lambda^{+}}<1/f_{\max}$. We here introduce two algorithms for searching this threshold:
  • Bisection Search: Given that $\mathcal{F}^{\lambda+\theta}$ is non-decreasing in $\theta$, we can apply a bisection search to gradually converge on the threshold $\theta^{\text{tr}}$. This method is a classical approach for locating thresholds in monotonic functions.
  • Intersection Search [58, Algorithm 2]: To improve the efficiency of solving constrained MDPs, [58] further exploits the piece-wise linear and concave (PWLC) structure Lagrangian cost function under finite state and action spaces. Unlike traditional bisection methods that locate the zero-crossing of a single monotonic function, intersection search identifies the intersection point of two tangents to the Lagrangian curve, thereby accelerating the search for the optimal Lagrange multiplier $\theta^{\text{tr}}$ with fewer iterations. We refer readers to [58, Section V.A] for a detailed discussion of this algorithm.
Once this threshold value is obtained, setting $\theta_{\lambda}^{\star}=\theta^{\text{tr}}$ ensures that the KKT conditions (52a)–(52c) are satisfied. What remains is to determine the optimal policy $\phi_{\lambda+\theta^{\star}_{\lambda}}$ that guarantees the KKT condition (52d). The structure of this optimal policy is provided in the following theorem.
@@
(ii). If $\mathcal{F}^{\lambda^{+}}\geq 1/f_{\max}$, the optimal policy is a stationary deterministic policy $\phi^{\star}_{\lambda}$ determined in (IV-B1);
(iiii). If $\mathcal{F}^{\lambda^{+}}<1/f_{\max}$ and $\mathcal{F}^{(\lambda+\theta^{\star}_{\lambda})^{+}}=1/f_{\max}$, the optimal policy is a stationary deterministic policy $\phi^{\star}_{(\lambda+\theta^{\star}_{\lambda})}$;
(iiiiii). If $\mathcal{F}^{\lambda^{+}}<1/f_{\max}$ and $\mathcal{F}^{(\lambda+\theta^{\star}_{\lambda})^{+}}>1/f_{\max}$, the optimal policy is a random mixture of two stationary deterministic policy $\phi^{\star}_{(\lambda+\theta^{\star}_{\lambda})^{+}}$ and $\phi^{\star}_{(\lambda+\theta^{\star}_{\lambda})^{-}}$:
@@
[…]